# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### 29 Feb 2024 v4

@@ I INTRODUCTION
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17]. In contrast, real time projection-based methods aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@ II Related Works
II-A Lidar Classification
@@
II-B Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14]. A common approach uses an active perception formulation to maximize entropy reduction [19, 10]. Ref. [12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs. Ref. [9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [22] similarly learn an information aware policy offline that proposes waypoints for the robot. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings for people.
II-C World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Prediction involves recognizing features common to tunnel networks to improve exploration strategies. Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. The planner uses this to explore rooms fully before continuing global exploration. This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time. Ref. [31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [8] instead learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations. This article describe a novel method for world prediction from 2D scans.
@@ III-E2 Viewpoint Selection
  • Subtract path distance from robot position to viewpoint

### 01 Mar 2024 v5

@@ I INTRODUCTION
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17]. Real time projection-based methods instead aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@ II Related Works
II-A Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14]. A common approach uses an active perception formulation to maximize entropy reduction [19, 10]. Ref. [12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs. Ref. [9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [22] similarly learn an information aware policy offline that proposes waypoints for the robot. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings during a search and rescue scenario.
II-B World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Prediction involves recognizing features common to tunnel networks to improve exploration strategies. Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. The planner uses this to explore rooms fully before continuing global exploration. This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time. Ref. [31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [8] learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations.
II-C Lidar Classification
@@ III-E2 Viewpoint Selection
  • Penalize distance from robot position to viewpoint